# DiFFPO: Training Diffusion LLMs to Reason Fast and Furious via Reinforcement Learning

### v1

@@ Abstract
We propose DiFFPO, Diffusion Fast and Furious Policy Optimization, a unified framework for training masked diffusion large language models (dLLMs) to reason not only better (furious), but also faster via reinforcement learning (RL). We first unify the existing baseline approach such as d1 (Zhao et al., 2025) by proposing to train surrogate policies via off-policy RL, whose likelihood is much more tractable as an approximation to the true dLLM policy. This naturally motivates a more accurate and informative two-stage likelihood approximation combined with importance sampling correction, which leads to generalized RL algorithms with better sample efficiency and superior task performance. Second, we propose a new direction of joint training efficient samplers/controllers of dLLMs policy. Via RL, we incentivize dLLMs’ natural multi-token prediction capabilities by letting the model learn to adaptively allocate an inference threshold for each prompt. By jointly training the sampler, we yield better accuracies with lower number of function evaluations (NFEs) compared to training the model only, obtaining the best performance in improving the Pareto frontier of the inference-time compute of dLLMs. We showcase the effectiveness of our pipeline by training open source large diffusion language models over benchmark math and planning tasks.
@@ 1 Introduction
Reinforcement Learning from Verifiable Rewards (RLVR, Lambert et al. (2024)) has achieved remarkable success in enhancing the reasoning capabilities of Large Language Models (LLMs) (Jaech et al., 2024; Guo et al., 2025). The fine-tuned Large Reasoning Models (LRMs) show drastic improvement in solving tasks which require strong reasoning capabilities, such as Math and Coding (Le et al., 2022; Shao et al., 2024). These LRMs match, and even surpass the performance of the best human players in math contests (Chen et al., 2025). Despite these successes, LRMs are notorious for long inference-time, and overthinking for easy questions (Chen et al., 2024; Su et al., 2025), which limit their applicability to scenarios that have less tolerance on latencies or inference budgets. An active line of recent research on efficient reasoning (Sui et al., 2025) proposed either training-free early stopping mechanisms, or generation-length-aware RL (Reinforcement Learning) objectives (Kim et al., 2025; Xu et al., 2025). However, the resulting RL fine-tuned models are still fundamentally bottle-necked by the left-to-right autoregressive (AR) decoding in decoder-only transformers (Vaswani et al., 2017), which will inevitably suffer from quadratic inference costs with respect to the length of reasoning traces.
Diffusion LLMs (dLLMs) (Lou et al., 2023; Sahoo et al., 2024; Shi et al., 2024; Nie et al., 2025; Ye et al., 2025), an emerging family of LLMs based on discrete-space diffusion models (Hoogeboom et al., 2021; Austin et al., 2021), have the natural premise of going beyond left-to-right generations to any-order generations and multi-token predictions. Properiatary dLLMs like Mercury (Khanna et al., 2025), Gemini Diffusion and Seed Diffusion (Song et al., 2025) maintain comparable quality to the state-of-the-art AR models while achieving up to 10 times of token throughput. In contrast with the extensive literature on post-training AR models, the relevant study for post-training dLLMs through RL in the literature so far remains rather limited. Whether RL can be used to enhance the reasoning capabilities of dLLMs still remains largely open, and how to design RL algorithms catered for dLLMs to incentives the capability of base model remains unexplored yet of crucial importance to improve frontier dLLMs.
@@
(1) We first propose an efficient RL post-training paradigm for fine-tuning dLLMs from an off-policy RL perspective. Concretely, we generalize recently proposed algorithms like d1 (Zhao et al., 2025) by proposing to train efficient surrogate policies with more tractable likelihood, as opposed to directly training base dLLMs policies, which requires extensive GPU memory to estimate and is inefficient to compute dLLM’s true likelihood. Based on our framework, we propose a new RL algorithm utilizing new surrogate policies by conditioning on additional latents at the response-generation level for better approximation, instead of only conditioning on prompts as in d1. Inspired by classical off-policy RL, we also introduce an importance-sampling correction term to address the distribution mismatch between the surrogate policies and the true dLLMs behavior policies. We provide both theoretical guarantees and strong empirical evidence, achieving evident performance gains over the baseline method, especially for planning tasks as in Figure 3.
@@
(2) Secondly, we innovatively propose to adopt efficient dLLMs samplers in RL post-training. Unlike prior work which post-trains the models based on a fixed (often inefficient) sampler, we directly train the model upon efficiently dLLMs samplers to inventive the base model to reason better and faster. To avoid overfitting to the utilized samplers, we propose to train a prompt-aware inference threshold motivated by the Entropy-Bounded (EB) sampler proposed in Ben-Hamu et al. (2025) via RL as opposed to a fixed threshold across different prompts. This approach naturally leverages the structural property of masked dLLMs for multi-token predictions, and lets the model perform inference based on a predicted “inference threshold” for the prompts. We enable efficient joint RL train of the model and sampler by treating the inference threshold as an additional token, which is thus compatible with our earlier proposed RL framework. We showcase that jointly training the model and sampler can yield improved post-trained models with better performance/accuracy under the same or even less compute than training the model only, unlocking a novel new direction for further research to investigate mechanisms between model and samplers in RL post training.
We term our pipeline as DiFFPO - Diffusion Fast and Furious Policy Optimization, for training masked dLLMs to reason not only better (furious), but also faster via RL. We demonstrate the effectiveness of DiFFPO in enhancing the inference-time performance of diffusion LLMs on benchmarking math and planning tasks, and clearly push the boundaries of designing scalable RL algorithms towards training efficient and capable LRMs.
The rest of the paper is organized as follows: In Section 2, we review the preliminaries of masked dLLMs, dLLMs samplers, and RL for discrete diffusion models. Section 3 presents our detailed algorithm of DiFFPO, with experimental results provided in Section 4. In Section 5 we discuss other relevant references and conclude with Section 6.
@@ 2 Preliminaries
In this work, we focus on masked discrete diffusion language models for simplicity, which has been shown to achieve the best performance among different formulations of dLLMs. Our framework is general and can be modified to other discrete diffusion models as well. For an introduction to dLLMs, we follow the presentations in MDLM (Sahoo et al., 2024).
@@
Since $\bm{x}$} is unknown during the forward process, MDLMs learn a function approximation ${\bm{x}}_{\theta}({\bm{z}}_{t},t)$, and approximate the true posterior with $p_{\theta}\left({\bm{z}}_{s}\mid{\bm{z}}_{t}\right)$ by replacing $\bm{x}$} with approximated ${\bm{x}}_{\theta}$ in Equation 2 when ${\bm{z}}_{t}=\bm{m}$.
dLLMs efficient samplers. For masked dLLMs, the backward process needs to choose among unmasked positions which to unmask first. This leaves an extra degree of freedom for designing efficient dLLMs samplers. Most existing works focus on a fixed number of multi-token predictions, such as random sampling or Top-$k$ sampling (Nie et al., 2025) by choosing the position based on some pre-defined scores. The scores are generally chosen from proxy metrics like confidence $s_{l}=\max_{v\in\mathcal{V}}p_{l}(v)$ or negative entropy $\mathcal{H}(p_{l})$ of the predictive distribution of dLLMs $p_{l}\in\Delta^{|\mathcal{V}|-1}$ on the position $l$. These top-$k$ based samplers are shown to be able to outperform randomly choosing $k$ positions (Kim et al., 2025).
In this work, our sampling scheme mainly follows the Entropy-Bounded (EB) sampler proposed in Ben-Hamu et al. (2025), which yields better accuracy and efficiency tradeoff than the Top-$k$ sampler that sweeps over different $k$’s. When jointly unmasking a sequence of variables $X=({\bm{x}}^{1},\cdots,{\bm{x}}^{L})$, the EB-sampler computes the entropy of the predictive distribution of dLLMs at each unmasked position as a cost of that position $c(l)=H(p^{\theta}({\bm{x}}^{l}\mid X))$, and then unmask all positions in $U$ with the maximum cardinality: $\sum_{l\in U}c(\ell)\leq\gamma$, where $\gamma$ is a predetermined threshold hyperparameter. If $\min_{l\in U}c(\ell)>\gamma$, EB sampler will only unmask the position with the smallest cost. Intuitively, small $\gamma$ will lead to behaviours similar to top-1 sampling, while large $\gamma$ will allow unmasking of several positions at the same time. Different $\gamma$ yields different samplers, which forms an inference-time compute frontier for dLLMs.
@@ 3 DiFFPO
3.1 Improved techniques in efficient model post-training
We first consider a fixed top-$k$ confidence score sampler, which is how the completions of prompts/questions are generated. We use $k=1$ for simplicity here to derive the loss objectives.
@@ 3.1 Improved techniques in efficient model post-training
(c) Visualization of off-policy RL for dLLM through optimizing the surrogate policy
Figure 2: Generation sample and average KL divergence error at each decoding timestep (100 prompts per dataset).
Rethinking via off-policy RL. Given the crude approximations utilized in d1 and the lack of any theoretical guarantee, we propose to formulate the RL task for dLLMs from an off-policy RL perspective by noticing the unique structure of the dLLMs. Since the true likelihood of dLLMs is inefficient to compute, our aim is to bypass it by: (1) find a surrogate policy that is close to the actual dLLM policy, while its conditional likelihood is tractable to compute; (2) optimize the surrogate policy via off-policy RL. A visualization of our pipeline can be found in Figure 2(c). Before we discuss how we choose the surrogate policies, we highlight that our current pipeline could provide a worst-case performance guarantee on the obtained fine-tuned models:
@@
The proof is a straightforward application of Pinsker’s inequality. Yet, a direct corollary of Theorem 1 implies that, the worst-case of the performance of $\pi^{\theta}$ can still be guaranteed if $\mathbb{E}_{o\sim\hat{\pi}_{\theta}(\cdot\mid c)}r(c,o)$ is large (e.g. after RL optimization) and two policies are close. Back to the dLLM setup, if we take a policy $\hat{\pi}_{\theta}$ which generates tokens on every position by conditioning only on the prompt, then this is the exact policy d1 utilized for likelihood approximation, which we have shown to be too coarse with a large KL-divergence. Thus we first construct better surrogate policies to make it closer to the true dLLM policy.
@@ 3.2 Unlocking better adaptive multi-token prediction via sampler post-training
Unlike the EB-sampler (Ben-Hamu et al., 2025) that uses a fixed threshold to decide which token(s) to unmask for all prompts, we propose to use a parameterized function to encode prompt features and output a predicted “inference threshold”. We obtain the dLLMs embeddings of the prompts by averaging over the hidden dimension features $f^{l}_{\theta}(c)$ of each position $l$, and train a linear mapping $\bm{w}$ on top of the embeddings (we omit the bias term here for clarity). We also introduce an upper threshold $\gamma_{\max}$, and map each feature to the inference threshold by
@@
We use $\pi^{w}_{\theta}$ to represent the sampler threshold policy which is dependent on both model weights $\theta$ and header weights $w$. To joint optimize the model and sampler, we apply a trick by treating the predicted threshold as an additional token to unmask at time step 0. Thus, given the group completion $(\gamma_{i},o_{i})$, we have yield our final DiFFPO loss as:
$\mathcal{J}_{\text{DiffPO}}=\mathcal{J}_{\text{DiffPO-model}}+\mathbb{E}_{c\sim\mathcal{D}}\sum_{i=1}^{G}\frac{1}{\left|o_{i}\right|}\min\left(\frac{\pi^{w}_{\theta_{-}}\left(\gamma_{i}\mid c\right)}{\pi^{w_{\text{old}}}_{\theta_{-}}\left(\gamma_{i}\mid c\right)}A_{i},\operatorname{clip}(\frac{\pi^{w}_{\theta_{-}}\left(\gamma_{i}\mid c\right)}{\pi^{w_{\text{old}}}_{\theta_{-}}\left(\gamma_{i}\mid c\right)},1-\varepsilon,1+\varepsilon)A_{i}\right),$ (15)
@@ 4 Experiments
We evaluate all the methods on math and planning benchmark tasks including GSM8K, Math, Sudoku, and Countdown. We use LoRA (Hu et al., 2022) with rank $r=128$ for parameter-efficient training – the same hyparameter utilized as in d1 (Zhao et al., 2025), along with other model generation setups for a fair comparison. In addition, we use the same five random reeds for the baseline d1 and our proposed algorithm, and report the best performing model among different seeds and checkpoints. For the dLLM inference setup, our experiments are all based on block size 32 and maximum sequence length of 256.
Improved sample efficiency and task performance. We first showcase the effectiveness of our algorithm when training a DiFFPO model with Equation 12 using a fixed top-$k$ sampler on math and planning benchmarks.
@@
Figure 3: Benchmark results ($\mathbin{\color[rgb]{0,0,1}{-}}$: DiFFPO; $\mathbin{\color[rgb]{1,0,0}{-}}$: DiFFPO without IS; $\mathbin{\color[rgb]{1,.5,0}{-}}$: d1 baseline).
We plot the reward as the training progresses in Figure 3. Our algorithm shows a clear margin over d1 in all tasks, especially in planning tasks. Concretely, using two-times mean-field approximation greatly improves the sample efficiency, and applying importance sampling correction further enhances the performance of RL training.
@@
Table 1: Benchmark results with statistics (Accuracy, $\Delta$, Effective Tokens (ETs)) under Top $k$ sampler. $\Delta$ indicates the accuracy improvement over LLaDA-8B-Instruct; positives are highlighted in green.
@@
We report concrete benchmark statistics in Table 1. Evidently, our proposed DiFFPO, combining both two times mean-field approximation and importance sampling, yields the most significant performance gain, which both components contribute to. We also compare the performance difference for the same model checkpoint utilizing different maximum length as in Table 2 (in which we denote ’2MF’ as an abbreviation for Two Times Mean-Field Approximation, and ’IS’ for importance sampling). We found that our DiFFPO trained models stably perform the best.
Training DiFFPO with efficient samplers. We also investigate the performance of RL-training on existing efficient samplers to gain the best performance. We tested EB samplers (Ben-Hamu et al., 2025), which yields a greater trade-off between accuracy and NFEs compared to the top-$k$ sampler that we used above. We found that the inference threshold $\gamma=0.1$ yields the best results in achieving the highest reward, on par with smaller $\gamma$’s and top-$k$ (Table 1, $k=2$), and better than larger $\gamma$’s which will typically yield a smaller average reward even after RL training.
@@
For a more detailed comparison, we draw the inference-time compute frontier for Countdown for two settings: 1) post-trained DiFFPO model with a fixed top-$k$ sampler (Equation 12); and 2) the model and sampler jointly trained by DiFFPO (Equation 15), under the same 256 maximum generation length in Figure 4. The dots are obtained by utilizing different base inference threshold in EB sampler on two models. The clear advantage of DiFFPO trained models (i.e. with sampler joint training) showcases the importance of choosing proper performant samplers when training the models via RL which is usually overloooked, and the effectiveness of our joint training mechanism in enhancing the inference-time compute frontier.

### v2

@@ Abstract
We propose DiFFPO, Diffusion Fast and Furious Policy Optimization, a unified framework for training masked diffusion large language models (dLLMs) to reason not only better (furious), but also faster via reinforcement learning (RL). We first unify the existing baseline approach such as d1 (Zhao et al., 2025) by proposing to train surrogate policies via off-policy RL, whose likelihood is much more tractable as an approximation to the true dLLM policy. This naturally motivates a more accurate and informative two-stage likelihood approximation combined with importance sampling correction, which leads to generalized RL algorithms with better sample efficiency and superior task performance. Second, we propose a new direction of joint training efficient samplers/controllers of dLLMs policy. Using RL, we incentivize dLLMs’ natural multi-token prediction capabilities by letting the model learn to adaptively allocate an inference threshold for each prompt. By jointly training the sampler, we yield better accuracies with lower number of function evaluations (NFEs) compared to training the model only, obtaining the best performance in improving the Pareto frontier of the inference-time compute of dLLMs. We showcase the effectiveness of our pipeline by training open source large diffusion language models over benchmark math and planning tasks.
@@ 1 Introduction
Reinforcement Learning from Verifiable Rewards (RLVR, Lambert et al. (2024)) has achieved remarkable success in enhancing the reasoning capabilities of Large Language Models (LLMs) (Jaech et al., 2024; Guo et al., 2025). The fine-tuned Large Reasoning Models (LRMs) show drastic improvement in solving tasks which require strong reasoning capabilities, such as Math and Coding (Le et al., 2022; Shao et al., 2024). These LRMs match, and even surpass the performance of the best human players in math contests (Chen et al., 2025). Despite these successes, LRMs are notorious for long inference-time, and overthinking for easy questions (Chen et al., 2024; Su et al., 2025), which limit their applicability to scenarios with low tolerance for latency or inference budgets. An active line of recent research on efficient reasoning (Sui et al., 2025) proposed either training-free early stopping mechanisms, or generation-length-aware RL (Reinforcement Learning) objectives (Kim et al., 2025; Xu et al., 2025). However, the resulting RL fine-tuned models are still fundamentally bottle-necked by the left-to-right autoregressive (AR) decoding in decoder-only transformers (Vaswani et al., 2017), which will inevitably suffer from quadratic inference costs with respect to the length of reasoning traces.
Diffusion LLMs (dLLMs) (Lou et al., 2023; Sahoo et al., 2024; Shi et al., 2024; Nie et al., 2025; Ye et al., 2025), an emerging family of LLMs based on discrete-space diffusion models (Hoogeboom et al., 2021; Austin et al., 2021), have the natural premise of going beyond left-to-right generations to any-order generations and multi-token predictions. Proprietary dLLMs like Mercury (Khanna et al., 2025), Gemini Diffusion and Seed Diffusion (Song et al., 2025) maintain comparable quality to the state-of-the-art AR models while achieving up to 10 times of token throughput. In contrast with the extensive literature on post-training AR models, the relevant study for post-training dLLMs through RL in the literature so far remains rather limited. Whether RL can be used to enhance the reasoning capabilities of dLLMs still remains largely open, and how to design RL algorithms catered for dLLMs to incentivize the capability of base model remains unexplored yet of crucial importance to improve frontier dLLMs.
@@
(1) We first propose an efficient RL post-training paradigm for fine-tuning dLLMs from an off-policy RL perspective. Concretely, we generalize recently proposed algorithms like d1 (Zhao et al., 2025) by proposing to train efficient surrogate policies with more tractable likelihood, as opposed to directly training base dLLMs policies, which requires extensive GPU memory to estimate and is inefficient to compute dLLM’s true likelihood. Based on our framework, we propose a new RL algorithm utilizing new surrogate policies by conditioning on additional latents at the response-generation level for better approximation, instead of only conditioning on prompts as in d1. Inspired by classical off-policy RL, we also introduce an importance-sampling correction term to address the distribution mismatch between the surrogate policies and the true dLLMs behavior policies. We provide both theoretical guarantees and strong empirical evidence, achieving evident performance gains over the baseline method, especially for planning tasks (like Countdown, Sudoku) in Figure 1.
@@
(2) Secondly, we innovatively propose to adopt efficient dLLMs samplers in RL post-training. Unlike prior work which post-trains the models based on a fixed (often inefficient) sampler, we directly train the model upon efficiently dLLMs samplers to inventive the base model to reason better and faster. To avoid overfitting to the utilized samplers, we propose to train a prompt-aware inference threshold motivated by the Entropy-Bounded (EB) sampler proposed in Ben-Hamu et al. (2025) via RL as opposed to a fixed threshold across different prompts. This approach naturally leverages the structural property of masked dLLMs for multi-token predictions, and lets the model perform inference based on a predicted “inference threshold” for the prompts. We enable efficient joint RL training of the model and sampler by treating the inference threshold as an additional token, which is thus compatible with our earlier proposed RL framework. We showcase that jointly training the model and sampler can yield improved post-trained models with better performance/accuracy under the same or even less compute than training the model only, unlocking a novel direction for further research to investigate mechanisms between model and samplers in RL post training.
We term our pipeline as DiFFPO, Diffusion Fast and Furious Policy Optimization, for training masked dLLMs to reason not only better (furious), but also faster via RL. We demonstrate the effectiveness of DiFFPO in enhancing the inference-time performance of diffusion LLMs on benchmarking math and planning tasks, and clearly push the boundaries of designing scalable RL algorithms towards training efficient and capable LRMs. The rest of the paper is organized as follows: In Section 2, we review the preliminaries of masked dLLMs, samplers, and RL for discrete diffusion models. Section 3 presents our detailed algorithm of DiFFPO, with experimental results provided in Section 4. In Section 5 we discuss other relevant references and conclude with Section 6.
@@ 2 Preliminaries
In this work, we focus on masked discrete diffusion language models for simplicity, which has been shown to achieve the best performance among different formulations of dLLMs. Our framework is general and can be adapted to other form of discrete diffusion models as well. For an introduction to dLLMs, we follow the presentations in MDLM (Sahoo et al., 2024).
@@
Since $\bm{x}$} is unknown, MDLMs learn a function approximation ${\bm{x}}_{\theta}({\bm{z}}_{t},t)$, and approximate the true posterior with $p_{\theta}\left({\bm{z}}_{s}\mid{\bm{z}}_{t}\right)$ by replacing $\bm{x}$} with approximated ${\bm{x}}_{\theta}$ in Equation 2 when ${\bm{z}}_{t}=\bm{m}$.
dLLMs efficient samplers. For masked dLLMs, the backward process needs to choose among unmasked positions which to unmask first. This leaves an extra degree of freedom for designing efficient dLLMs samplers. Most existing works focus on a fixed number of multi-token predictions, such as random sampling or Top-$k$ sampling (Nie et al., 2025) by choosing the position based on some pre-defined scores. The scores are generally chosen from proxy metrics like confidence $s_{l}=\max_{v\in\mathcal{V}}p_{l}(v)$ or negative entropy $\mathcal{H}(p_{l})$ of the predictive distribution of dLLMs $p_{l}\in\Delta^{|\mathcal{V}|-1}$ on the position $l$. These Top-$k$ based samplers are shown to outperform randomly choosing $k$ positions (Kim et al., 2025). Above mechanism is also referred as ‘remasking’ in (Nie et al., 2025) if we think of the equivalent procedure by unmasking every position and remasking those with low scores.
In this work, our sampling scheme mainly follows the Entropy-Bounded (EB) sampler proposed in Ben-Hamu et al. (2025), which yields better accuracy and efficiency tradeoff than the Top-$k$ sampler that sweeps over different $k$’s. When jointly unmasking a sequence of variables $X=({\bm{x}}^{1},\cdots,{\bm{x}}^{L})$, the EB-sampler computes the entropy of the predictive distribution of dLLMs at each unmasked position as a cost of that position $c(l)=H(p^{\theta}({\bm{x}}^{l}\mid X))$, and then unmask all positions in $U$ with the maximum cardinality: $\sum_{l\in U}c(\ell)\leq\gamma$, where $\gamma$ is a predetermined threshold hyperparameter. If $\min_{l\in U}c(\ell)>\gamma$, EB sampler will only unmask the position with the smallest cost. Intuitively, small $\gamma$ will lead to behaviours similar to Top-1 sampling, while large $\gamma$ will allow unmasking of several positions at the same time. Different $\gamma$ forms an inference-time compute frontier for dLLMs.
@@ 3 DiFFPO
3.1 Improved techniques in sample efficient model post-training
For the sampler of the dLLM, we first consider a fixed Top-$k$ confidence based remasking, which is how the completions of prompts/questions are generated. We use $k=1$ for simplicity here to derive the loss objectives without the loss of generality.
@@ 3.1 Improved techniques in sample efficient model post-training
(c) Comparison of 2-MF approximation with baseline mean field approximation.
Figure 3: Visualization of off-policy RL for dLLM through optimizing the surrogate policy
Rethinking via off-policy RL. Given the crude approximations utilized in d1 and the lack of any theoretical guarantee, we propose to formulate the RL task for dLLMs from an off-policy RL perspective by noticing the unique structure of the dLLMs. Since the true likelihood of dLLMs is inefficient to compute, our aim is to bypass it by: (1) find a surrogate policy that is close to the actual dLLM policy, while its conditional likelihood is tractable to compute; (2) optimize the surrogate policy via off-policy RL. A visualization of our pipeline can be found in Figure 3: we assume that the direction of $r$ ($\hat{r}$) encourages the true dLLM (surrogate) policy to generate higher average reward separately. Then if we optimize the surrogate policy following $\hat{r}$, the performances of true dLLM policy under the same model weights $\theta_{t}$ gets improved if the projections of it to the linear space $\theta_{0}+r$, i.e. $\theta_{t}^{\perp}$, move along the direction of $r$.
Before we discuss how we choose the surrogate policies, we highlight that our current pipeline could provide a worst-case performance guarantee on the obtained fine-tuned models:
@@
The proof is a straightforward application of Pinsker’s inequality. Yet, a direct corollary of Theorem 1 implies that, the worst-case of the performance of $\pi^{\theta}$ can still be guaranteed if $\mathbb{E}_{o\sim\hat{\pi}_{\theta}(\cdot\mid c)}r(c,o)$ is large (e.g. after RL optimization) and two policies are close. Back to the dLLM setup, if we take a policy $\hat{\pi}_{\theta}$ which generates tokens on every position by conditioning only on the prompt, then this is the exact policy d1 utilized for likelihood approximation, which we have shown to be too coarse with a large KL-divergence. Thus we are motivated to construct better surrogate policies to make it closer to the true dLLM policy.
@@
In Figure 2(c), we show the clear empirical evidence on GSM8K dataset that two times mean field approximation indeed yields better approximation to the true likelihood in terms of reducing the KL divergence error. We provide similar results for other datasets in Appendix A.1.
@@ 3.2 Unlocking better adaptive multi-token prediction via sampler post-training
Prompt adaptive threshold. Unlike the EB-sampler (Ben-Hamu et al., 2025) that uses a fixed threshold to decide which token(s) to unmask for all prompts, we propose to use a parameterized function to encode prompt features and output a predicted “inference threshold”. We obtain the dLLMs embeddings of the prompts by averaging over the hidden dimension features $f^{l}_{\theta}(c)$ of each position $l$, and train a linear mapping $\bm{w}$ on top of the embeddings (we omit the bias term here for clarity). We also introduce an upper threshold $\gamma_{\max}$, and map each feature to the inference threshold by
@@
Joint Training of the model and sampler. We use $\pi^{w}_{\theta}$ to represent the sampler threshold policy which is dependent on both model weights $\theta$ and header weights $w$. To joint optimize the model and sampler, we apply a trick by treating the predicted threshold as an additional token to unmask at time step 0. Thus, given the group completion $(\gamma_{i},o_{i})$, we have yield our final DiFFPO loss as:
$\mathcal{J}_{\text{DiffPO}}=\mathcal{J}_{\text{DiffPO-model}}+\mathbb{E}_{c\sim\mathcal{D}}\sum_{i=1}^{G}\frac{1}{\left|o_{i}\right|}\min\left(\frac{\pi^{w}_{\theta_{-}}\left(\gamma_{i}\mid c\right)}{\pi^{w_{\text{old}}}_{\theta_{-}}\left(\gamma_{i}\mid c\right)}A_{i},\operatorname{clip}(\frac{\pi^{w}_{\theta_{-}}\left(\gamma_{i}\mid c\right)}{\pi^{w_{\text{old}}}_{\theta_{-}}\left(\gamma_{i}\mid c\right)},1-\varepsilon,1+\varepsilon)A_{i}\right).$ (15)
@@ 4 Experiments
We evaluate all the methods on math and planning benchmark tasks including GSM8K, Math, Sudoku, and Countdown. We use LoRA (Hu et al., 2022) with rank $r=128$ for parameter-efficient training and also adopt the 4-bit quantized version of the model for efficiency, same as d1. Our training is conducted on 8 NVIDIA-A100 40GB GPUs, batch size of 3 per GPU and use gradient accumulation steps of 4. We grid search the best learning rate and clip ratio for each method, and adopt the same all other hyparameters utilized in d1 (Zhao et al., 2025) like AdamW optimizer for a fair comparison. In addition, we use the same five random reeds for the baseline d1 and our proposed algorithm, and report the best performing model among different seeds and checkpoints.
For the dLLM inference setup, our experiments are all based on block size 32 and maximum sequence length of 256. We adopt a Top-$k$ confidence based remasking mechanism for inference when training the model only without training the sampler. We choose $k=2$, the same setup as d1, resulting the effective diffusion steps being 128.
Improved sample efficiency and task performance. We first showcase the effectiveness of our algorithm when training a DiFFPO model with Equation 12 on math and planning benchmarks.
@@
Figure 4: Benchmark results ($\mathbin{\color[rgb]{0,0,1}{-}}$: d1; $\mathbin{\color[rgb]{0.0235,0.5412,0.2118}{-}}$: wd1; $\mathbin{\color[rgb]{1,.5,0}{-}}$: ours).
We plot the reward as the training progresses in Figure 4. Our algorithm shows a clear margin over d1 in all tasks, especially in planning tasks like Countdown and Sudoku. Our algorithm also outperforms concurrent work like wd1 (Tang et al., 2025) that also targets at improving d1 on the same reasoning datasets.  [^3] Note that since our paper’s main focus is on the probability ratio and wd1 focuses on advantage function, it is possible to combine these two methods. We leave the investigation as future work.
@@
Table 1: Benchmark results with statistics (Accuracy, $\Delta$, Effective Tokens (ETs)) under Top $k$ sampler. $\Delta$ indicates the accuracy improvement over LLaDA-8B-Instruct; positives are highlighted in green. * Results follow from evaluation in Dream 7B (Ye et al., 2025).
@@
We report concrete benchmark statistics in Table 1. The row of ’+ Importance Sampling’ corresponds to adopting two times mean-field and importance sampling at the same time, while the row of ’+ Two Times Mean-Field’ corresponds to adopting two times mean-field only. Evidently, our proposed DiFFPO, combining both two times mean-field approximation and importance sampling, yields the most significant performance gain, which both components contribute to.
We also compare the performance difference for the same model checkpoint utilizing different maximum length as in Table 2 (in which we denote ’2MF’ as an abbreviation for Two Times Mean-Field Approximation, and ’IS’ for importance sampling). We found that our DiFFPO trained models stably perform the best.
Training DiFFPO with efficient samplers. We also investigate the performance of RL-training on existing efficient samplers to gain the best performance. We tested upon EB samplers (Ben-Hamu et al., 2025), which yields a greater trade-off between accuracy and NFEs compared to the top-$k$ sampler that we used above. We found that the inference threshold $\gamma=0.1$ yields the best results in achieving the highest reward, on par with smaller $\gamma$’s and top-$k$ (Table 1, $k=2$), and better than larger $\gamma$’s which will typically yield a smaller average reward even after RL training.
@@
For a more detailed comparison, we draw the inference-time compute frontier for Countdown for two settings: 1) post-trained DiFFPO model with a fixed top-$k$ sampler (Equation 12); and 2) the model and sampler jointly trained by DiFFPO (Equation 15), under the same 256 maximum generation length in Figure 5. The dots are obtained by utilizing different base inference threshold (here we choose [0.05, 0.1,0.2, 0.5,1,5]) in EB sampler on two models. The clear advantage of DiFFPO trained models (i.e. with sampler joint training) showcases the importance of choosing proper performant samplers when training the models via RL which is usually overloooked, and the effectiveness of our joint training mechanism in enhancing the inference-time compute frontier.